OpenStudy (anonymous):

help please x+7y=39 3x-2y=2 solve using substitution

OpenStudy (anonymous):

@iGreen can u help me please

OpenStudy (igreen):

Subtract 7y to both sides: x + 7y = 39 x = -7y + 39 Distribute -7y + 39 for 'x' in the 2nd equation: 3x - 2y = 2 3(-7y + 39) - 2y = 2 Distribute 3 into the parenthesis: -21y + 117 - 2y = 2 Combine like terms, what's -21y - 2y?

OpenStudy (anonymous):

-23?

OpenStudy (igreen):

-23y, yes. So we have: -23y + 117 = 2 Subtract 117 to both sides, what's 2 - 117?

OpenStudy (anonymous):

-115? can u please tell me how u got 117 in the first place

OpenStudy (igreen):

I distributed 3 into the parenthesis.. 3 * -7y = -21y 3 * 39 = 117

OpenStudy (igreen):

Anyway, we have: -23y = -115 Divide -23 to both sides, what's -115 / -23?

OpenStudy (anonymous):

oh okay I understand now.. y=5?

OpenStudy (igreen):

Yep, you got it.. Now we plug it into any of the two equations to get the 'x-value'. x + 7y = 39 x + 7(5) = 39 Simplify & Solve for 'x'

OpenStudy (anonymous):

a little stuck got x=38?

OpenStudy (igreen):

No..what's 7 * 5?

OpenStudy (anonymous):

nevermind I had 3 at the end and not 39... 7*5=35 I got that part

OpenStudy (igreen):

Yes, so we have: x + 35 = 39 Subtract 35 to both sides

OpenStudy (anonymous):

so the final answer is (4,5)?

OpenStudy (igreen):

Yep, you got it.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

x+7y=39 x=39-7y------(1) 3x-2y=2 3(39-7y)-2y=2 117-21y-2y=2 117-23y=2 117-2=23y 115=23y y=115\23 y=5 3x-2(5)=2 3x-10=2 3x=12 x=4

OpenStudy (anonymous):

yup that's what I got thanks @sangineni